what is the value of x in the system shown ?
x+y=7
3x-y=5

can you explain it in details?

Answer by addingup(3677)

Add them together:
x+y=7
+
3x-y=5
-------
4x=12 (+y and -y cancel each other) Now divide both sides by 4:
x= 3
In the first equation:
3+y=7 Subtract 3 on both sides:
y= 4
Let's plug both these numbers in the second equation:
3(3)-4= 5 This is correct. Now the first equation:
3+4=7 This is also correct. We have the correct answers.